Let f(x) =x²-a. Show that the Newton-raphson method leads to the recurrent

1
Expert's answer
2021-11-09T16:29:44-0500

A recurrence relation is an equation that expresses each element of a sequence as a function of the preceding ones.


by the Newton-Raphson method:


"x_{n+1}=x_n-\frac{f(x_n)}{f'(x_n)}"


"f'(x)=2x"


recurrence relation:


"x_{n+1}=x_n-\frac{x_n^2-a}{2x_n}"